Rolodex Inc. is in the process of determining its capital budget for the next fiscal year. The firms current capital structure, which it considers to be optimal, is contained in the following balance sheet: Note: For this problem, use the book value of the items to get the capital structure. However, you normally want to use the market values. Long-term debt is the only debt capital structure account. Add common stock, capital in excess of par, and RE to get common equity. Rolodex Inc. Balance Sheet (in Millions of Dollars) Current assets $105 Fixed assets Total assets $360 Accounts payable Other current liabilities Long-term debt Preferred stock Common stock (15 million shares at par) Contributed capital in excess of par Retained earnings Total liabilities and equity 35 25 90 60 15 30 105 $360 255 Discussions between the firms financial officers and the firms investment and commercial bankers have yielded the following information: Rolodex can borrow $30 million from its bank at a pretax cost of 11 percent.Rolodex can borrow $30 million from its bank at a pretax cost of 11 percent. Rolodex can borrow $25 million by issuing bonds at a net price of $875 per bond. The bonds would carry a 13 percent coupon rate and mature in 15 years. Additional debt can be issued at a 16.5 percent pretax cost Preferred stock can be issued at a pretax cost of 18 percent. Rolodex expects to generate $120 million in net income and pay $2.15 per share in dividends Hint: RE = Net income-total common dividends·There are 15m shares shown as outstanding in the balance sheet. The $2.15 per share dividend (D1) represents a growth of 5.5 percent over the previous years dividend. This growth rate is expected to continue for the foreseeable future. The firms stock currently is trading at $16 per share. Rolodex can raise external equity by selling common stock at a net price of $14.50 per share. Rolodexs marginal tax rate is 40 percent. Round your answers to two decimal places. 10.12% would be entered as 10.12 What is the weight on debt? . What is the after-tax cost of debt for the second class of debt? . What is the internal cost of equity? 0What is the internal cost of equity? What is the external cost of equity? Does debt or equity have the lowest breakpoint? Enter debt or equity. Compute Rolodexs marginal cost of capital schedule. Weighted marginal cost of capital First increment: Second increment: Third increment: Additional funds:
Answer #1

(a)Total debt on balance sheet = 90 Million

Total market value of equity = 15 Million * $16/share = 240 Million

Total preferred Stock = 60 Million

Total capital = 390 Million

Weight of debt = 90/390 = 0.2308 = 23.08%

(b) Pretax cost of debt for second class of debt =RATE(nper,pmt,pv,fv) where nper =15,pmt=13%*1000 = 130, pv=875 and fv =1000

Pretax cost of debt for second class of debt = RATE(15,130,-875,1000) = 15.15%

After tax cost of debt for the second class = 15.15*(1-tax rate) = 15.15*(1-0.4) = 9.09%

(c) Internal cost of equity = D1/P0 + g = 2.15/16 +0.055 = 18.94% (We use the current share price of $16 for internal equity)

(d) External cost of equity, we use the new issue share price of $14.50

External cost of equity = 2.15/14.50+0.055 = 0.2033 = 20.33%

Note:We have answered four sub parts. Please note that only four sub-parts can be answered at a time. Kindly post the others separately for experts to answer